OpenStudy (anonymous):

Find f(x) IF: f(x+y) = f(x) + f(y) - f(x)f(y) And f(0) is not 1 And f'(0) = -1

OpenStudy (anonymous):

@experimentX The basic fault there in your thing is that, x and y are INDEPENDENT variables. Therefore change in x does not change y in any way. Hence, While differenciating you can differenciate it w.r.t. x as f'(x+y) = f'(x) - f'(x)f(y). Now I put x=0 , => f'(y) + 1 = f(y) Thats how far I've gotten so far.

OpenStudy (anonymous):

dy/dx -y = -1

OpenStudy (experimentx):

that's a linear DE.

hartnn (hartnn):

let me tell u what i did f'(x) = lim y->0 f(x+y) - f(x) / y = lim y->0 f(y)(1-f(x)) / y put x=0 -1 = ...... since the limit exist , there must be common factor that will cancel out 1-f(0) cannot be y.it must be constant so f(y) = y f(x) = x

hartnn (hartnn):

is there any mistake ^^ ??

OpenStudy (anonymous):

hmm.

OpenStudy (anonymous):

that's pretty cool @hartnn

OpenStudy (anonymous):

Okay. Yeah. I see. Linear D.E.

OpenStudy (anonymous):

@hartnn Thats pretty cool but, I guess, f(y) = y When y tends to zero only. Otherwise we cannot say anything.

hartnn (hartnn):

yup, i also realised that f'(0) is not -1 then...so i m incorrect

OpenStudy (anonymous):

lim y->0 f(x+y) - f(x) / y wouldn't you need the limit def. of the partial though?

OpenStudy (anonymous):

f'(0) is -1 for the record.

OpenStudy (anonymous):

Nice. Thanks guys.

OpenStudy (experimentx):

1-e^(x+y) = (1 - e^x)+(1-e^y) - (1 - e^x)*(1-e^y) = 2 - (e^x+e^y) - 1 + (e^x+e^y) - e^(x+y) = 1 - e^(x+y)